# Immersed figure-8 annuli and anyons

### v2

@@ Abstract
Immersion (i.e., local embedding) is relevant to the physics of topologically ordered phases through entanglement bootstrap. An annulus can immerse in a disk or a sphere as a “figure-8”, which cannot be smoothly deformed to an embedded annulus. We investigate a simple problem: is there an Abelian state on the immersed figure-8 annulus, locally indistinguishable from the ground state of the background physical system? We show that if the answer is affirmative, a strong sense of isomorphism must hold: two homeomorphic immersed regions must have isomorphic information convex sets, even if they cannot smoothly deform to each other on the background physical system. We explain why to care about strong isomorphism in physical systems with anyons and give proof in the context of Abelian anyon theory. We further discuss a connection between immersed annuli and anyon transportation in the presence of topological defects. In appendices, we discuss related problems in broader contexts.
@@ 1 Introduction
To motivate the usage of immersed regions, let us first recall why embedded regions are of interest in the study of topologically ordered systems with emergent anyonic excitations. On closed manifolds, such as tori, there can be multiple locally indistinguishable ground states. The information that distinguishes those ground states cannot be seen on any ball-shaped subsystem. Meaningful differences can, nevertheless, be detected if we examine subsystems with nontrivial topologies, such as annuli and (punctured) tori.
@@
Immersed regions are a broader class of regions. They are locally embedded but need not be globally embedded. Because of this, states in immersed regions can use multiple copies of local Hilbert spaces of the original physical system. For instance, states on the figure-8 annulus in Fig. 1 use the Hilbert space of the overlapping region (the region covered by two sheets) twice. Local embedding is sufficient for an information convex set to be defined [16] because we only needed local consistency of state in the information convex set with the reference state. Similarly, information convex sets are shown to be isomorphic under smooth deformation of the immersed region. To our knowledge, the topological notion of immersion rarely appeared in physics before. A notable exception is a sequence of connections observed by Hastings [19] and with Freedman [20] on invertible phases and the classification of quantum cellular automata.
Figure 1: An immersed “figure-8” annulus ($X_{8}$) on a background 2-dimensional physical system on which we have a reference state $\sigma$. The background is often taken to be a ball or a sphere. Entanglement bootstrap axioms A0 and A1 are imposed on bounded radius balls such as the ball (yellow) illustrated. Entropy combinations $\Delta(B,C)$ and $\Delta(B,C,D)$ are defined in Eq. (1).
@@
Despite the progress mentioned above, the study of immersed regions in entanglement bootstrap is still in its early stages. Some topological manifolds can immerse in topologically distinct ways, even in a sphere. A simple example is the immersed “figure-8” annulus, which cannot be smoothly deformed to an embedded annulus. One natural question is whether two homeomorphic regions $\Omega,\Omega^{\prime}$ immersed in the $n$-dimensional sphere in inequivalent classes, still have isomorphic information convex sets (Conjecture 6 and 14), which we call the strong isomorphism conjecture. One may also wonder if smooth deformation (and also “tunneling” Sec. 3.1) between different immersed regions can give interesting automorphism of information convex sets. These two problems are closely related, as we shall explain, and both are open problems.
@@
A crucial intuition we used in the proof is that anyons transported along the figure-8 annulus cannot be permuted. We discuss a thought experiment “transportation experiment” (Sec. 5), that broadens the scope of this idea and applies it to physical contexts with topological defects. It relates the property of immersed annulus that thickens the transportation loop to the question of if anyons are permuted after the transportation. Many ideas and questions generalize to higher dimensions. In the appendix, we discuss a few tools useful in thinking about such generalizations.
@@ 2 Background
This section is a review of some necessary background information for entanglement bootstrap. We shall focus on the concept of immersed regions and related facts. In particular, we explain why there are precisely two classes of immersed annuli in a sphere; see Table 1.
@@ 2.1 Setup and axioms
We consider entanglement bootstrap setup in two spatial dimensions (2D). We consider a reference state $\sigma$ on a 2D surface. Unless stated otherwise, we consider a reference state on the sphere $\mathbf{S}^{2}$. (We used bold font to emphasize that a reference state is defined on the sphere.) In this case, $\sigma=|\psi\rangle\langle\psi|$. [^2] Starting with a ball or a sphere is equivalent due to the “completion trick” [16]. We shall take $\mathbf{S}^{2}$ to be explicit. The important thing is that we do not need a state on a topologically nontrivial manifold. We assume that the total Hilbert space is a tensor product of on-site Hilbert spaces from sites (or cells) of a coarse-grained lattice. The local Hilbert space for each coarse-grained site is finite-dimensional.
@@
Here $S_{A}|_{\rho}$ means $S(\rho_{A})=-{\rm Tr}(\rho_{A}\ln\rho_{A})$ is the von Neumann entropy. The strong subadditivity [22] implies useful inequalities of similar (but diverse) forms. A particularly useful one to keep in mind is that, for any $\rho_{ABCD}$, $\Delta(B,C,D)_{\rho}\geq I(A:C|B)_{\rho}\geq 0$. Here $I(A:C|B)\equiv S_{AB}+S_{BC}-S_{B}-S_{ABC}$ is the conditional mutual information.
@@
The axioms A0 and A1 stated that the entropy conditions hold precisely. This can be a pretty strong requirement even though whole classes of solvable models satisfy these conditions precisely [23]. Nevertheless, there is evidence that precisely satisfied axioms are too strong to accommodate chiral topological orders. [^3] At least, this should be true if the local Hilbert space on each site is finite-dimensional. One piece of evidence comes from an unpublished work with Daniel Ranard, Ting-Chun Lin, and Isaac Kim, where a commuting projector parent Hamiltonian is identified. From the physical standing point, the right way to appreciate these axioms is that we expect them to hold more accurately in larger systems. The first appearance of entropy in physics is in thermodynamics, where the prediction works best for large systems. We draw the parallel and argue that entanglement bootstrap predictions are also relevant for systems with approximate axioms. In fact, for chiral phases, a certain formula initially argued used precise area law [8] found confirmation with another theoretical approach [24] in the range of validity of the latter.
The axioms are imposed on an intermediate length scale, which is much smaller than the total system size and larger than the correlation length and (microscopic) lattice spacing. This can be viewed as one instance of starting from the “middle” [25] rather than the “bottom” [26], where entanglement bootstrap may be considered as one of those most quantum cases. Indeed, suppose the two axioms are satisfied on some intermediate scale. In that case, the same conditions hold on all larger length scales, and they define a renormalization group fixed point in this sense.
@@
It is a conjecture that if the violations of the axioms are small enough, coarse-grain the lattice further will result in more accurate axioms. These are “threshold conjectures”; see Chapter 11 of [27] for a version and an explanation of why a threshold $0<\delta<\ln 2$ of violation of A1 is consistent with current examples from spurious topological entanglement entropy, non-Abelian anyons, domain walls, and defects.
Although we are mainly interested in systems where the axioms hold, some concepts and insights developed in entanglement bootstrap, such as the idea of considering quantum states on immersed regions, may be useful in broader contexts as well.
2.2 Immersed annuli and its classification
Immersed regions are locally embedded regions. For each immersed region, an information convex set can be defined [14, 16]. An information convex set is a set of density matrices on a region $\Omega$ that can locally extend and are locally identical to the reference state $\sigma$. Any locally embedded region has a “good enough” set of balls to check such consistency conditions, and thus, an information convex set $\Sigma(\Omega)$ is defined. Embedded regions, also called “subsystems,” are special cases of immersed regions. (We shall denote “$\Omega$ is immersed in $Y$” by $\Omega\looparrowright Y$.)
The isomorphism theorem says if two immersed regions ($\Omega,\Omega^{\prime}\looparrowright\mathbf{S}^{2}$) can be smoothly [^4] In entanglement bootstrap, we have a coarse-grained lattice rather than a smooth manifold, in which context “smooth deformation” always means applying elementary steps; see Section 3.3 of [14]. deformed to each other, the information convex sets associated with them are isomorphic. We denote this as
@@ 2.2 Immersed annuli and its classification
[…]
@@
Below, we shall be interested in immersed annuli. As the isomorphism theorem implies, only the topological class is of interest. We explain below that there are only two topological classes of immersed annuli, denoted by $X$ and $X_{8}$; see Table 1. (We adopt two alternative ways to draw $X_{8}$. One is as in Fig. 1, where the overlap part is shown as two layers, one on top and it covers another layer. The second way is as Table 1, where the overlap part is transparent so that we see both layers. We note that the “top and bottom” in the first way is introduced for drawing purposes only, and this is not part of the data to specify immersion. We shall frequently use the first approach as it simplifies drawing.) We shall be interested in knowing if $\Sigma(X)\cong\Sigma(X_{8})$, which is not answered by the isomorphism theorem.

### v4

@@ Abstract
Immersion (i.e., local embedding) is relevant to the physics of topologically ordered phases through the entanglement bootstrap. An annulus can be immersed in a disk or a sphere as a “figure-8”, which cannot be smoothly deformed to an embedded annulus. We investigate a simple problem: is there an Abelian state on the immersed figure-8 annulus, locally indistinguishable from the ground state of the background physical system? We show that if the answer is affirmative, a strong sense of isomorphism must hold: two homeomorphic immersed regions must have isomorphic information convex sets, even if they cannot smoothly deform to each other on the background physical system. We explain why to care about strong isomorphism in physical systems with anyons and give proof in the context of Abelian anyon theory. We further discuss a connection between immersed annuli and anyon transportation in the presence of topological defects. In the appendices, we discuss related problems in broader contexts.
@@ 1 Introduction
To motivate the usage of immersed regions, let us first recall why embedded regions are of interest in the study of topologically ordered systems with emergent anyonic excitations. On closed manifolds, such as tori, there can be multiple locally indistinguishable ground states. The information that distinguishes those ground states cannot be seen on any ball-shaped subsystem. Meaningful differences can, nevertheless, be detected if we examine subsystems with interesting topologies, such as annuli and (punctured) tori.
@@
Immersed regions are a broader class of regions. They are regions locally embedded in the physical system but need not be globally embedded. Because of this, states in immersed regions can use multiple copies of local Hilbert spaces of the original physical system. For instance, the figure-8 annulus can be immersed in a ball (Fig. 1), and quantum states on the figure-8 annulus use the Hilbert space of the overlapping region (the region covered by two sheets) twice. Local embedding is sufficient for an information convex set to be defined [16, 17, 18] because we only needed local consistency of state in the information convex set with the reference state. Information convex sets are shown to be isomorphic under smooth deformation of the immersed region. These isomorphisms not only preserve the shape of the convex set but also preserve various distance measures between density matrices in the convex set.
To our knowledge, the topological notion of immersion has rarely appeared in physics before. A notable exception is a sequence of connections observed by Hastings [19] and with Freedman [20] on invertible phases and the classification of quantum cellular automata.
Figure 1: An immersed “figure-8” annulus ($X_{8}$) on a background 2-dimensional physical system. The background physical system, often taken to be a ball or a sphere, is equipped with a reference state $\sigma$. For the reference state, entanglement bootstrap axioms A0 and A1 are imposed on bounded radius balls such as the yellow ball illustrated. Entropy combinations $\Delta(B,C)$ and $\Delta(B,C,D)$ are defined in Eq. (1).
@@
Despite the progress mentioned above, the study of immersed regions in entanglement bootstrap is still in its early stages. Some topological manifolds can immerse in topologically distinct ways, even in a sphere. A simple example is the immersed “figure-8” annulus, which cannot be smoothly deformed to an embedded annulus. One natural question is whether two homeomorphic regions $\Omega,\Omega^{\prime}$ immersed in the $n$-dimensional sphere in inequivalent classes, still have isomorphic information convex sets (Conjecture 9 and 17), which we call the strong isomorphism conjecture. One may also wonder if smooth deformation (and also “tunneling” Sec. 3.1) between different immersed regions can give interesting automorphism of information convex sets. These two problems are closely related, as we shall explain, and both are open problems.
@@
A crucial intuition we used in the proof is that anyons transported along the figure-8 annulus cannot be permuted. We discuss a thought experiment “transportation experiment” (Sec. 5), that broadens the scope of this idea and applies it to physical contexts with topological defects. It relates the property of the immersed annulus that thickens the transportation loop to the question of whether anyons are permuted after the transportation. Many ideas and questions generalize to higher dimensions. In the appendix, we discuss a few useful tools for thinking about such generalizations.
@@ 2 Background
In this section, we review some necessary background for the entanglement bootstrap [14, 15, 16, 17]. We review the setup and axioms, and then review a central concept, the information convex set. We then focus on the concept of immersed regions and related facts. In particular, we explain why there are precisely two classes of immersed annuli in a sphere; see Table 1.
@@ 2.1 Setup and axioms
We consider the entanglement bootstrap setup in two spatial dimensions (2D). We consider a reference state $\sigma$ on a 2D surface $M$. Unless stated otherwise, we consider a reference state on the sphere $\mathbf{S}^{2}$, letting $M=\mathbf{S}^{2}$. (We used bold font to emphasize that a reference state is defined on the sphere.) In this case, $\sigma=|\psi\rangle\langle\psi|$. [^2] Starting with a ball or a sphere is equivalent due to the “completion trick” [16]. We shall take $\mathbf{S}^{2}$ to be explicit. The important thing is that we do not need a state on a topologically nontrivial manifold, such as a torus. We assume that the total Hilbert space is a tensor product of on-site Hilbert spaces from sites (or cells) of a coarse-grained lattice. The local Hilbert space for each coarse-grained site is finite-dimensional.
@@
Here $S_{A}|_{\rho}$ means $S(\rho_{A})=-{\rm Tr}(\rho_{A}\ln\rho_{A})$ is the von Neumann entropy. The strong subadditivity [22] implies useful inequalities of similar (but diverse) forms. A particularly useful one to keep in mind is that, for any $\rho_{ABCD}$, $\Delta(B,C,D)_{\rho}\geq I(A:C|B)_{\rho}\geq 0$. Here $I(A:C|B)\equiv S_{AB}+S_{BC}-S_{B}-S_{ABC}$ is the conditional mutual information. Whenever the labels of regions are listed together, we mean the union of disjoint regions; for instance, $AB$ refers to the union of non-overlapping regions $A$ and $B$, where $A$ and $B$ do not share any qudits.
@@
The axioms A0 and A1 stated that the entropy conditions hold precisely. This can be a pretty strong requirement even though there are classes of solvable models that satisfy these conditions precisely [23]. Nevertheless, there is evidence that precisely satisfied axioms are too strong to accommodate chiral topological orders, at least when the local Hilbert space on each site is finite-dimensional.333 One piece of evidence comes from [24], where a commuting projector parent Hamiltonian is identified. For a different argument that is based entirely on the quantum state, see [25]. From a physical standpoint, the right way to appreciate these axioms is that we expect them to hold more accurately in larger systems. The first appearance of entropy in physics is in thermodynamics, where the prediction works best for large systems. We draw the parallel and argue that entanglement bootstrap predictions are also relevant for systems with approximate axioms. In fact, for chiral phases, a certain formula initially argued used precise area law [8] found confirmation with another theoretical approach [26] in the range of validity of the latter.
The axioms are imposed on an intermediate length scale, which is much smaller than the total system size and larger than the correlation length and (microscopic) lattice spacing. This can be viewed as one instance of starting from the “middle” [27] rather than the “bottom” [28], where entanglement bootstrap may be considered as one of the most quantum cases. Indeed, suppose the two axioms are satisfied on some intermediate scale. In that case, the same conditions hold on all larger length scales, and they define a renormalization group fixed point in this sense.
@@
It is a conjecture that if the violations of the axioms are small enough, coarse-graining the lattice further will result in more accurate axioms. In other words, there may be a threshold of violation of the axioms, below which we expect the error of the axioms to decay with the length scale. See Chapter 11 of [29] for a version of the conjecture and an explanation of why a threshold $0<\delta<\ln 2$ of violation of A1 is consistent with known examples from non-Abelian anyons, domain walls, defects, and spurious topological entanglement entropy.
2.2 Information convex set
The information convex set is a central concept that is useful for understanding the main results of the paper. We will first recall its definition. Despite the effort given to the precise definition, we emphasize that knowing the precise definition is not required to understand the main result of the paper. An alternative route to read the rest of the paper is to treat the information convex set as a “black box”, whose property is given by the isomorphism theorem and structure theorems we discuss. Readers taking this alternative route could still understand the main technical innovation without much loss.
We say two density matrices $\rho_{AB}$ and $\lambda_{AB}$ are indistinguishable on subsystem $A$ if their reduced density matrices are identical on $A$, namely ${\rm Tr}_{B}\rho_{AB}={\rm Tr}_{B}\lambda_{AB}$. Alternatively, we say
[…]
which reads “the states $\rho_{AB}$ is consistent with $\lambda_{A}$,” where $\lambda_{A}={\rm Tr}_{B}\lambda_{AB}$ is the reduced density matrix.
The essence of information convex set ($\Sigma(\Omega)$) is a set of density matrices on region $\Omega$, such that any element is indistinguishable from the reference state $\sigma$ (introduced in section 2.1) on a set of local regions. The formal definition also involves the ability to thicken the region $\Omega$, and thus it is a bit subtle. In fact, there are several alternative definitions, but we choose not to discuss all of them.
We first state the definition of the information convex set for an embedded region $\Omega$ (Definition 1). For the definition to be general, we denote the space manifold of the physical system as $M$. When the region $\Omega$ is embedded, we write $\Omega\hookrightarrow M$. We say an embedded region is a subsystem. We then discuss the more general definition that is applicable to immersed regions (Definition 3) following a formal definition of immersed region (Definition 2).
Definition 1 (Information convex set for embedded region [14]).
For an embedded region $\Omega$, which can be thickened to $\Omega_{+}$, the information convex set $\Sigma(\Omega)$, for a given reference state $\sigma$, is the set of density matrices
$\Sigma(\Omega)\equiv\{\rho_{\Omega}|\text{ conditions }1,\,2\},$ (3)
where the two conditions are
  1. 1.
    $\rho_{\Omega}={\rm Tr}_{\Omega_{+}\setminus\Omega}\,\rho_{\Omega_{+}}$, where $\rho_{\Omega_{+}}$ is a density matrix on $\Omega_{+}$.
  2. 2.
[…]
Figure 2: Regions used in the definition of information convex set. (a) $\Omega$ is an annulus embedded in the plane. $\Omega_{+}\supset\Omega$ is its thickening. $b\subset\Omega_{+}$ is a bounded radius disk. (b) An immersed region $\Omega$, which is homeomorphic to a punctured torus.
The embedded regions $\Omega,\Omega_{+}$ and $b$ that appeared in the definition are illustrated in Fig. 2(a). A few remarks are in order.
  1. 1.
    An embedded region can be thought of as a pair $\Omega=(\omega,\mathfrak{p})$ where $\omega$ is an abstract topological manifold of the same dimension as $M$ (possibly with boundaries). $\mathfrak{p}:\omega\to M$ is the embedding map, which maps $\omega$ to $M$, such that the image is homeomorphic to $\omega$ through the map $\mathfrak{p}$. We can thus identify the embedded region with the image, and say $\Omega\subset M$. While this explanation may seem optional (as embedded regions are intuitively well-understood), we find it useful in the comparison with the formal definition of immersed regions below (Definition 2). The information of the embedded region is less than the embedding map because composing the embedding map with automorphism $\mathfrak{a}$ of $\omega$ (that is $\mathfrak{p}^{\prime}=\mathfrak{p}\circ\mathfrak{a})$ defines the same embedded region.
  2. 2.
    The fact that the disk $b$ can be chosen from balls with a certain bounded radius is an important detail. It enabled the proof of various theorems in previous works [14, 15, 16]. However, omitting the requirement of boundedness of radius still gives a valid alternative definition.
  3. 3.
    It follows from this definition that $\Sigma(\Omega)$ must be a compact convex set.
Immersed regions: Informally, immersed regions are regions that are locally embedded in the background manifold $M$. We take the following definition of immersed regions. It is essentially the 2nd half of the definition of immersed regions in [16], with a minor simplification.
Definition 2 (Immersed region).
An immersed region is a pair $\Omega=({\omega},\mathfrak{p})$. Here $\omega$ be a topological manifold of the same dimension as the physical system $M$. $\mathfrak{p}:\omega\to M$ is an continuous map such that, for every point $x\in\omega$, there exists a disk-like neighborhood $n(x)$ such that when we restrict the domain of the map $\mathfrak{p}$ to $n(x)$, it gives an homeomorphism onto the image $b(\mathfrak{p}(x))=\mathfrak{p}(n(x))\subset M$, where $b(\mathfrak{p}(x))$ is a bounded radius disk.
We use $\Omega\looparrowright M$ to indicate that $\Omega$ is a region that is immersed in $M$. When we say $\Omega$ is an annulus, we mean the topology of $\omega$ in the pair $\Omega=(\omega,\mathfrak{p})$ is an annulus. We shall refer to the continuous map $\mathfrak{p}$ obeying the requirement in Definition 2 as the immersion map. For our purpose, replacing the immersion map by $\mathfrak{p}^{\prime}=\mathfrak{p}\circ\mathfrak{a}$ give the same immersed region, where $\mathfrak{a}$ is an automorphism of $\omega$.
Immersed regions can be specified by drawing a layered region ($\Omega$) onto the background manifold ($M$), such as the figure-8 annulus $X_{8}$ in Fig. 1, and the punctured torus $\Omega$ in Fig. 2(b). Note that such figures can determine the immersion map up to the automorphism (of $\omega$), which is precisely what we need. Embedded region is a special type of immersed region. The advantage that immersed regions bring us is that we can sometimes use the qudits in the original physical system more than once. We recycle the qudit!
The precise meaning of recycling qubits is that we can define a pullback Hilbert space on the immersed region $\Omega=(\omega,\mathfrak{p})$. Suppose the original physical system has the total Hilbert space as a tensor product of onsite Hilbert spaces. We can find a set of sites on $\omega$ as the preimages of the physical sites under $\mathfrak{p}$. The Hilbert space on the immersed region $\Omega$ is defined to be the tensor product of the local Hilbert spaces on these sites. In general, the Hilbert space is larger than that associated with the image $\mathfrak{p}(\omega)\subset M$. The subsets of the image that are covered twice in Fig. 2(b), each contributes two copies of the qudits. We also have the ability to pull back pieces of the reference state on a neighborhood $n(x)\subset\omega$. We shall denote such pull-back reference state as $\sigma_{n}^{[\mathfrak{p}]}$. Note, however, that we cannot directly pull back a “global” reference state on a nontrivially immersed $\Omega$; this is in contrast with what we can do for embedded regions.
Information convex set for immersed regions: Because the previous definition of an information convex set (Definition 1) only requires comparing the density matrices on a set of bounded radius balls (neighborhood of points), it is reasonable to extend the definition to immersed regions. Indeed, this is possible.
Definition 3 (Information convex set for immersed region [16]).
For an immersed region $\Omega\looparrowright M$, which can be thickened into immersed region $\Omega_{+}=(\omega_{+},\mathfrak{p})$, the information convex set $\Sigma(\Omega)$, for a given reference state $\sigma$, is the set of density matrices
$\Sigma(\Omega)\equiv\{\rho_{\Omega}|\text{ conditions }1,\,2\},$ (4)
where the two conditions are
  1. 1.
    $\rho_{\Omega}={\rm Tr}_{\Omega_{+}\setminus\Omega}\,\rho_{\Omega_{+}}$, where $\rho_{\Omega_{+}}$ is a density matrix on $\Omega_{+}$.
  2. 2.
[…]
In practice, we may use the drawings of $\Omega\looparrowright M$ and $b\subset M$ [as in Fig. 2(b)] to describe the second condition, i.e. the local consistency condition. Such a description omits the explicit mentioning of neighborhood $n$ and the immersion map $\mathfrak{p}$; this does require pointing to the layer in the drawing of $\Omega$ that $b$ talks to.
2.3 Isomorphism, immersed annuli and its classification
As explained in the previous section, immersed regions are regions locally embedded in the background physical system. For an immersed region $\Omega$, an information convex set $\Sigma(\Omega)$ can be defined (Definition 3). Embedded regions, also called “subsystems,” are special cases of immersed regions. We shall denote “$\Omega$ is immersed in $M$” by $\Omega\looparrowright M$. We are mainly interested in the case where the background physical system $M$ is the sphere.
We say two embedded regions $\Omega$ and $\Omega^{\prime}$ are related by an elementary step [14] if we can partition the two regions as $\Omega=AB$ and $\Omega^{\prime}=ABC$ in a relation shown in Fig. 3. Here $BCD$ is a partition of a bounded radius disk that we impose axiom A1. Here, more precisely, we call $\Omega\to\Omega^{\prime}$ an elementary step of extension and call $\Omega^{\prime}\to\Omega$ an elementary step of restriction. The definition generalizes when the regions $\Omega$ and $\Omega^{\prime}$ are immersed.
Figure 3: An illustration related to elementary steps. $A$ can be large and has an arbitrary topology, and it can be immersed. $BCD$ is a partition of a bounded radius (embedded) disk that we impose axiom A1.
[…]
There is an important theorem, called the isomorphism theorem [14, 16], in entanglement bootstrap, under axioms A0 and A1. The isomorphism theorem says for $\Omega,\Omega^{\prime}\looparrowright\mathbf{S}^{2}$,
@@ 2.3 Isomorphism, immersed annuli and its classification
By isomorphism $\cong$”, we mean the existence of an quantum channel $\Phi:\Sigma(\Omega)\to\Sigma(\Omega^{\prime})$, which can be reversed by another channel $\widetilde{\Phi}:\Sigma(\Omega^{\prime})\to\Sigma(\Omega)$. Such channels are completely determined by the sequence of elementary steps. As an isomorphism, some nontrivial structures of information convex sets are preserved, following from the existence of such channels:
  • (i)
    The geometry of the convex set is preserved. Namely, the extreme points of $\Sigma(\Omega)$ and $\Sigma(\Omega^{\prime})$ are mapped to each other, and the map commutes with the convex combination:
    $\Phi(p\rho+(1-p)\lambda)=p\,\Phi(\rho)+(1-p)\Phi(\lambda),\quad\quad\forall\rho,\lambda\in\Sigma(\Omega),\quad p\in[0,1].$ (6)
  • (ii)
    The distance measure between any two elements is preserved. For instance, the trace distance $D(\cdot,\cdot)$ and the fidelity $F(\cdot,\cdot)$ satisfy:
    $D(\rho,\lambda)=D(\Phi(\rho),\Phi(\lambda)),\quad F(\rho,\lambda)=F(\Phi(\rho),\Phi(\lambda)),\quad\forall\rho,\lambda\in\Sigma(\Omega).$ (7)
  • (iii)
    The von Neumann entropy difference between any two elements is preserved:
    $S(\rho)-S(\lambda)=S(\Phi(\rho))-S(\Phi(\lambda)),\quad\forall\rho,\lambda\in\Sigma(\Omega).$ (8)
Classification of immersed annuli: As the isomorphism theorem suggests, what is interesting is the topological classification of immersed regions. Below, we shall be interested in immersed annuli. We explain below that there are only two topological classes of immersed annuli, denoted by $X$ and $X_{8}$; see Table 1. (We adopt two alternative ways to draw $X_{8}$. One is as in Fig. 1, where the overlap part is shown as two layers, one on top and covering another layer. The second way is as Table 1, where the overlap part is transparent so that we see both layers. We note that the “top and bottom” in the first way is introduced for drawing purposes only, and this is not part of the data to specify immersion. We shall frequently use the first approach as it simplifies drawing.) We shall be interested in knowing if $\Sigma(X)\cong\Sigma(X_{8})$, which is not answered by the isomorphism theorem.